OpenStudy (anonymous):

Simplify the expression ten x minus 4 plus 3 x.

OpenStudy (anonymous):

10.x-4+3.x=13.x-4

OpenStudy (jim766):

10x -4 + 3x = 13x - 4 first step is to combine like terms on ea side of = 13x - 4 = 13x - 4 get the variable alone with a coefficient of 1 -13x -13x -4 = -4 if you end up eliminating the variable, there are 2 possibilities: 1. you end up with a true statement infinite solutions 2. you end up with a false statement no solution

OpenStudy (anonymous):

7x -4